# Floor Plan-Guided Visual Navigation Incorporating Depth and Directional Cues

### v2

@@ Introduction
Implementation
In our implementation, the multi-head attention consists of four heads, each with four layers. We choose Square Cosine Noise Schedule (nichol2021improved) to train the diffusion policy. We use the AdamW optimizer, with the learning rate controlled by PyTorch’s CosineAnnealingLR scheduler. The maximum learning rate is set to $0.0001$, and the maximum number of iterations is $20$. We set $\alpha$ to $0.1$. The training of GlocDiff is performed on $8$ NVIDIA $4090$ GPUs, with a batch size of $64$ per GPU. GlocDiff is trained for $N=20$ epochs and reaches convergence.
Experiments
To evaluate our approach thoroughly, we conduct extensive experiments within the simulation environment, including baseline comparisons and ablation studies. Furthermore, we deploy our policy on a real AGV, demonstrating its practical utility, robustness, and generalization capabilities in real-world settings.
Main Experiments
Table 1: The details of the baselines with respective localization and planning modules, as well as training data size.
@@
Loc-A (GT) GT Pose A -
Loc-FloDiff (F3) (li2024flona) F3Loc (chen2024f3loc) Loc-FloDiff 180180M
Loc-FloDiff (GT) (li2024flona) GT Pose Loc-FloDiff 180180M
@@
Setup
We conduct our experiments in the Gibson (xia2018gibson) simulator. The testing episodes are sourced from FloNa (li2024flona), comprising fifty unseen scenes, with each scene containing ten episodes. To ensure safe navigation, we check whether the proposed action would cause a collision before execution. If so, we intervene by instructing the agent to perform a 45° left turn and re-plan a new action sequence. The navigation is deemed successful if the final agent position falls within $\tau_{d}$ meters of the goal and the total number of interventions does not exceed $\tau_{c}$.
In our experiments, we select six baselines, with their specific components outlined in Tab. 1. Both Loc-FloDiff (F3) and GlocDiff (F3) employ F3Loc (chen2024f3loc), an SOTA end-to-end floor plan localization model, as their localization module. Both Loc-FloDiff (GT) and GlocDiff (GT) use the ground truth (GT) pose as localization results. The Loc-A (GT) leverages ground-truth localization and directly follows the shortest path computed by the A algorithm on the floor plan. In GlocDiff (VO), a VO algorithm (freda2025pySLAM) is used to track the agent’s pose, serving as the localization module. Due to the lack of scale information, we manually set the scale to 0.040.04m, based on the average inter-frame displacement in our dataset.
Unlike the dense sampling strategy used for collecting data when training Loc-FloDiff, we sample every five frames to ensure that each observation in a trajectory is sampled only once. As a result, GlocDiff is trained on $1.2$ million samples, whereas Loc-FloDiff is trained on $180$ million samples. For evaluation, we adopt Success Rate (SR) and Success Weighted by Path Length (SPL) as our metrics. We evaluate all baselines using the distance threshold values ($\tau_{d}$) of $0.25$, $0.30$, and $0.35$, as well as the intervention count values ($\tau_{c}$) of $10$, $30$, $50$, and $\infty$.
Table 2: Performance of different baselines under various thresholds. Bold numbers indicate the best results.
Method Loc-FloDiff (F3) GlocDiff (F3) GlocDiff (VO) Loc-A (GT) Loc-FloDiff (GT) GlocDiff (GT)
$\tau_{d}(m)$ $\tau_{d}(m)$ $\tau_{d}(m)$ $\tau_{d}(m)$ $\tau_{d}(m)$ $\tau_{d}(m)$
$0.25$ $0.30$ $0.35$ $0.25$ $0.30$ $0.35$ $0.25$ $0.30$ $0.35$ $0.25$ $0.30$ $0.35$ $0.25$ $0.30$ $0.35$ $0.25$ $0.30$ $0.35$
SR(%) $\uparrow$ $\tau_{c}$ $10$ $4.80$ $6.00$ $6.40$ $53.40$ $56.00$ $57.80$ $68.60$ $69.80$ $70.20$ $16.60$ $25.40$ $38.60$ $39.00$ $44.20$ $44.60$ $79.60$ $80.40$ $80.60$
$30$ $11.60$ $14.00$ $15.00$ $69.20$ $70.00$ $71.80$ $80.80$ $82.20$ $82.80$ $18.20$ $28.40$ $43.80$ $53.40$ $60.80$ $61.00$ $92.80$ $92.80$ $93.20$
$50$ $17.80$ $20.60$ $22.40$ $77.80$ $78.80$ $80.60$ $86.60$ $87.60$ $87.80$ $18.50$ $29.40$ $45.60$ $59.20$ $68.40$ $68.40$ $94.20$ $94.20$ $94.20$
$\infty$ $26.80$ $30.80$ $33.00$ $86.40$ $87.40$ $89.40$ $89.80$ $90.40$ $90.60$ $18.60$ $29.60$ $45.80$ $66.00$ $75.80$ $76.00$ $97.60$ $97.60$ $97.60$
[…]
$30$ $6.40$ $8.25$ $8.98$ $49.35$ $52.35$ $54.16$ $59.07$ $61.91$ $63.59$ $18.10$ $28.30$ $43.51$ $44.10$ $50.40$ $51.20$ $70.38$ $72.57$ $74.06$
$50$ $8.07$ $9.98$ $10.80$ $52.61$ $55.77$ $57.55$ $61.31$ $63.90$ $65.50$ $18.50$ $29.30$ $45.11$ $46.60$ $53.40$ $54.10$ $70.98$ $72.98$ $74.48$
$\infty$ $9.58$ $11.70$ $12.50$ $54.77$ $57.94$ $59.75$ $62.57$ $65.04$ $66.64$ $18.50$ $29.40$ $45.31$ $48.20$ $55.20$ $56.00$ $72.36$ $74.36$ $75.86$
Figure 4: Trajectory visualization. The visualization of trajectories for four baselines across six episodes.
Quantitative Results
Tab. 2 presents the quantitative results on the FloNa test set. We discuss three key observations:
  • GlocDiff demonstrates a consistent and substantial performance improvement over both Loc-FloDiff and Loc-A in terms of SR and SPL, irrespective of whether ground-truth or estimated poses are employed. This performance gap becomes increasingly pronounced under stricter thresholds of $\tau_{d}$ and $\tau_{c}$. Such notable gains underscore GlocDiff’s ability to effectively exploit the complementary nature of floor plan priors and visual observations, thereby markedly enhancing the agent’s capacity for obstacle avoidance and efficient path planning.
  • When the localization results shift from GT poses to predictions from the F3Loc module, Loc-FloDiff experiences a significant decline in performance, with the SR dropping in the range of $34.2\%$ to $47.8\%$ and the SPL dropping in the range of $32.38\%$ to $43.60\%$. In contrast, the performance of GlocDiff shows a more moderate decline, with the SR falling by $28.20\%$ to $39.80\%$ and the SPL decreasing by $28.56\%$ to $32.68\%$. This highlights that our model exhibits greater robustness to localization noise.
  • The SR of GlocDiff (VO) is approximately $50\%$ higher than that of GlocDiff (F3), indicating that visual odometry (VO) is a more suitable localization module for floor plan navigation compared to F3Loc. Since VO can continuously track the agent’s pose, it remains relatively stable for indoor navigation tasks, despite cumulative errors. In contrast, F3Loc maps observations to poses, introducing room-level errors, which are unacceptable for floor plan navigation. We quantitatively evaluate the localization accuracy of both VO and F3Loc to support this claim; please refer to the supplementary material for details. Moreover, as the $\tau_{c}$ increases, the SPL gap between GlocDiff (VO) and GlocDiff (GT) gradually narrows, indicating the accumulated error of VO does not significantly affect the planning ability.
Qualitative Results
Fig. 4 visualizes the trajectories of different models across six episodes.
Ablation Study
Ablation Study on Depth and Directional Cues
We further evaluate the effects of different guidance on the policy. To assess the impact of the local depth guidance, we use an EfficientNet to directly encode the RGB observations, obtaining RGB context to replace the original depth context, while keeping other design aspects consistent with those of GlocDiff. The resulting ablation model is denoted as w/o depth. To analyze the influence of the directional guidance, we directly utilize the target point to replace the shortest path, while also keeping other aspects consistent with GlocDiff. This modification results in another ablation model, referred to as w/o path. The two models described above are both trained for $20$ epochs and reach convergence. We evaluate SR, SPL, and the mean Number of Collisions (NoC) of two models using GT poses.
Table 3: Ablation results. Bold and underlined numbers indicate the best and second-best results, respectively.
Model w/o depth w/o path GlocDiff (GT)
$\tau_{d}(m)$ $\tau_{d}(m)$ $\tau_{d}(m)$
$0.25$ $0.30$ $0.35$ $0.40$ $0.25$ $0.30$ $0.35$ $0.40$ $0.25$ $0.30$ $0.35$ $0.40$
SR(%) $\uparrow$ $\tau_{c}$ $10$ $64.20$ $65.00$ $66.40$ $66.80$ $70.60$ $71.40$ $72.40$ $72.80$ $79.60$ $80.40$ $80.60$ $80.80$
$30$ $81.00$ $81.60$ $82.80$ $83.00$ $82.00$ $82.40$ $83.00$ $83.20$ $92.80$ $92.80$ $93.20$ $93.20$
$50$ $85.60$ $87.00$ $88.00$ $88.20$ $85.00$ $85.60$ $86.00$ $86.20$ $94.20$ $94.20$ $94.20$ $94.20$
$\infty$ $92.20$ $93.00$ $93.40$ $93.40$ $89.00$ $89.40$ $89.60$ $89.80$ $97.60$ $97.60$ $97.60$ $97.60$
SPL(%) $\uparrow$ $\tau_{c}$ $10$ $47.22$ $49.83$ $52.20$ $53.60$ $50.44$ $52.74$ $54.84$ $55.90$ $62.62$ $64.85$ $66.44$ $67.71$
$30$ $55.98$ $58.70$ $60.98$ $62.40$ $55.90$ $58.07$ $60.09$ $61.21$ $70.38$ $72.57$ $74.06$ $75.34$
$50$ $58.07$ $60.84$ $63.11$ $64.58$ $57.24$ $59.49$ $61.42$ $62.54$ $70.98$ $72.98$ $74.48$ $75.76$
$\infty$ $60.15$ $62.93$ $64.90$ $66.30$ $58.51$ $60.71$ $62.53$ $63.66$ $72.36$ $74.36$ $75.86$ $77.16$
NoC $\downarrow$ $\tau_{c}$ $10$ $4.46$ $4.34$ $4.23$ $4.17$ $3.70$ $3.63$ $3.52$ $3.49$ $3.09$ $2.99$ $2.95$ $2.89$
$30$ $9.46$ $9.19$ $8.84$ $8.71$ $7.87$ $7.77$ $7.54$ $7.46$ $5.38$ $5.16$ $5.10$ $5.01$
$50$ $12.66$ $12.19$ $11.63$ $11.45$ $11.06$ $10.96$ $10.63$ $10.51$ $6.65$ $6.39$ $6.34$ $6.24$
$\infty$ $21.78$ $20.41$ $19.61$ $19.42$ $31.93$ $31.82$ $31.06$ $30.41$ $10.81$ $10.55$ $10.49$ $10.39$
Figure 5: Performance of different baselines on the test episodes in Mobridge (left) and Spotswodd (right). The left column illustrates the traversed trajectory, while the right column showcases the diverse actions generated by different models based on the same observation.
Figure 6: Three real-world test episodes. The left columns present the top-down view of successful episodes, with $1$ and 55 indicating the start and goal positions. The right columns show the AGV’s first-person RGB observations at the corresponding positions.
As outlined in Tab. 3, GlocDiff (GT) achieves the best performance across all cases. Three key conclusions can be drawn. First, the removal of either the local guidance or the global guidance significantly degrades the model’s performance. This indicates that both global and local guidance contribute to enhancing the robot’s ability to navigate safely in complex environments. Specifically, the global guidance provides essential support for efficient navigation across the floor plan, while the local guidance improves the understanding of the scene’s geometric structure, thereby aiding in the identification of traversable regions. Second, the performance of the w/o depth model is lower than that of the w/o path model under low $\tau_{c}$ thresholds in terms of SR and SPL, but the trend reverses as the threshold increases. This indicates that removing depth features impairs the agent’s ability to avoid obstacles, which in turn leads to more frequent interventions and reduced performance at lower $\tau_{c}$ thresholds. However, as the intervention threshold increases, planning ability becomes more critical, leading to a performance drop for the w/o path model due to the lack of global planning guidance. Third, when $\tau_{c}$ approaches infinity, the w/o path model surprisingly incurs more collisions than the w/o depth variant, which appears counterintuitive. We hypothesize that the absence of shortest path guidance compels the agent to follow unnecessarily prolonged and circuitous trajectories, thereby increasing the likelihood of collisions.
We select two episodes, each from a different unseen scene, and visualize the results in Fig. 5. When the agent navigates to the same goal, different models generate distinct actions due to their varying capabilities. The w/o depth model encounters collisions with walls and furniture in both cases, indicating the role of local guidance in generating collision-free navigation actions. In Mobridge, the w/o path model generates incorrect actions at the intersection due to the lack of guidance from the optimal path.
Table 4: Performance comparison of w/o perturbation and GlocDiff under two levels of noise.
Model w/o perturbation GlocDiff
$\tau_{d}(m)$ $\tau_{d}(m)$
$0.25$ $0.35$ $0.25$ $0.35$
SR(%) $\uparrow$ $u_{1}(m)$ $\tau_{c}$ $10$ $61.00$ $69.00$ $73.00$ $75.00$
$30$ $73.00$ $75.00$ $86.00$ $88.00$
$50$ $80.00$ $82.00$ $92.00$ $93.00$
$50$ $86.00$ $90.00$ $96.00$ $96.00$
$u_{2}(m)$ $\tau_{c}$ $10$ $42.00$ $45.00$ $65.00$ $65.00$
$30$ $52.00$ $57.00$ $69.00$ $72.00$
$50$ $60.00$ $65.00$ $77.00$ $83.00$
$\infty$ $77.00$ $78.00$ $86.00$ $87.00$
SPL(%) $\uparrow$ $u_{1}(m)$ $\tau_{c}$ $10$ $42.08$ $46.38$ $53.95$ $57.39$
$30$ $57.15$ $60.80$ $61.86$ $65.61$
$50$ $61.79$ $62.24$ $64.84$ $68.79$
$\infty$ $63.41$ $65.82$ $65.63$ $69.58$
$u_{2}(m)$ $\tau_{c}$ $10$ $29.90$ $33.38$ $44.40$ $44.40$
$30$ $34.41$ $39.93$ $52.07$ $56.77$
$50$ $37.73$ $40.84$ $57.23$ $59.21$
$\infty$ $41.61$ $46.19$ $61.14$ $63.75$
Ablation Study on GT Perturbation
We retrain our model with the GT perturbation trick removed and refer to this revised version as w/o perturbation. We then assess the robustness of GlocDiff and w/o perturbation to localization noise across ten unseen environments. Specifically, we introduce two levels of noise, $u_{1}(m)\sim\mathcal{U}(0.0,0.5)$ and $u_{2}(m)\sim\mathcal{U}(1.0,2.0)$, to the GT localization results and evaluate the performance of GlocDiff and w/o perturbation under both conditions. As shown in Tab. 4, GlocDiff outperforms w/o perturbation across all conditions, and the performance of w/o perturbation degrades more severely than that of GlocDiff as the noise level increases. This highlights that incorporating noise during training significantly enhances the model’s robustness to localization errors.
Real World Deployment
This section presents the results of deploying the GlocDiff on an AGV in a real apartment without any fine-tuning. The AGV is equipped with a mobile base, an RGB camera, and an NVIDIA Jetson AGX Orin for model execution. We scan the apartment using Scaniverse and apply post-processing to generate a floor plan with characteristics consistent with those in the dataset. For reliable localization, we employ the AGV’s single-line LiDAR-based odometry for pose estimation, with the robot’s initial pose pre-calibrated to ensure stable performance. On the Orin device, the model achieves an inference speed of approximately $1.21\text{Hz}$. The actions generated by GlocDiff are converted into linear and angular velocities via a PD controller.
We evaluate the system on $10$ episodes, with $\tau_{c}=0$ and $\tau_{d}=20\text{cm}$. Each pair is tested three times to ensure reliability, resulting in a final success rate of $40\%$. Fig. 6 showcases three navigation episodes, highlighting the generalization of obstacle avoidance and planning capabilities to unseen real-world environments. Please refer to the attached demo for a more intuitive illustration.
Conclusion
In summary, we propose an efficient and practical visual navigation framework based on floor plans, leveraging local and global condition extraction alongside a diffusion policy. Our method addresses the challenge of efficiently integrating visual information with the floor plan for optimal path planning. Meanwhile, visual odometry ensures stable localization, enhancing real-world applicability. Looking forward, we foresee the integration of multi-modal data, such as language or object information, further improving adaptability and robustness, paving the way for intelligent indoor service robots in complex environments.
Limitations
Despite the promising results, our approach has certain limitations. First, we assume a consistent scale between the floor plan and the real-world environment. However, in real scenarios, floor plans may exhibit distortions or scale mismatches, which could undermine navigation accuracy. Second, although our diffusion-based policy improves the agent’s obstacle avoidance ability, it still cannot guarantee fully collision-free behavior. To enable reliable deployment in practical applications, this issue must be further addressed.
Figure A1: Failure case. The left column presents the GT path and the generated path on the floor plan. The right column showcases an RGB observation along with the predicted depth and visualizes the GT path, the generated path, and the shortest path on the observation.
Table A1: Error Comparison of Localization Methods
@@ Conclusion
Appendix A Failure Case
Fig. A1 illustrates a typical failure case where our model generates a path (red line) that deviates from the optimal path (green line), despite the global guidance (blue line) being correct. This occurs because, from current viewpoint, the region suggested by global guidance exhibits low depth. Consequently, GlocDiff produces a detour toward a region with higher depth to avoid a perceived collision risk. However, this perceived low depth is merely an illusion caused by the narrow field of view (FOV) and does not reflect that the suggested area is truly impassable. This indicates that a larger FOV could help generate more accurate actions.
Appendix B Performance of VO and F3Loc
We additionally evaluate the accuracy of two localization modules, i.e. VO (freda2025pySLAM) and F3Loc (chen2024f3loc). The input images for localization are of size $512\times 512\times 3$, with both the vertical and horizontal fields of view set to $90$ degrees. We employ the multi-frame inference mode of the pre-trained F3Loc model to estimate camera poses. For VO, we initialize the camera pose of the first frame using the ground truth. Due to the lack of an absolute scale, we manually specify the initial displacement between the first and second frames as $0.04$ m, corresponding to the average inter-frame motion observed in our dataset. We randomly sample a total of fifty episodes from ten scenes and compute the mean and median errors between the predicted and ground-truth poses at each time step. As shown in Tab. A1, VO demonstrates a higher average accuracy compared to F3Loc. Moreover, F3Loc exhibits occasional outliers, suggesting that its estimated poses abruptly jump to distant locations within the environment. Such discontinuous pose estimations are intolerable for GlocDiff. In contrast, although VO is susceptible to cumulative drift, the noise introduced remains within a range that GlocDiff can robustly tolerate, making VO a more suitable choice in practice.
Table A2: Performance Comparison of Different Prediction Horizons on GlocDiff (GT).
Steps
$16$ $32$ $48$
SR(%) $\uparrow$ $\tau_{c}$ $10$ $77.40$ $79.60$ $79.00$
$30$ $89.80$ $92.80$ $91.80$
$50$ $94.00$ $94.20$ $94.40$
$\infty$ $97.80$ $97.60$ $97.40$
SPL(%) $\uparrow$ $\tau_{c}$ $10$ $62.69$ $66.60$ $65.00$
$30$ $67.83$ $70.38$ $69.09$
$50$ $68.62$ $70.98$ $70.32$
$\infty$ $71.89$ $72.36$ $72.23$
Appendix C Ablation on Prediction Horizon

### v3

@@ III-E Training
III-F Implementation
In our implementation, the multi-head attention consists of four heads, each with four layers. We choose Square Cosine Noise Schedule [19] to train the diffusion policy. We use the AdamW optimizer, with the learning rate controlled by PyTorch’s CosineAnnealingLR scheduler. The maximum learning rate is set to $0.0001$, and the maximum number of iterations is $20$. We set $\alpha$ to $0.1$. The training of GlocDiff is performed on $8$ NVIDIA $4090$ GPUs, with a batch size of $64$ per GPU. GlocDiff is trained for $N=20$ epochs and reaches convergence.
IV Experiments
We evaluate our method through comprehensive simulation experiments (including baseline comparisons and ablation studies) and subsequently deploy the policy on a physical AGV.
TABLE II: Baseline Methods and Their Components.
@@ IV Experiments
Loc-A* (GT) GT Pose A* -
DD-PPO (GT) GT Pose $D$-PPO 2.52.5B
Map-A* (GT) GT Pose A* -
Loc-FloDiff (GT) GT Pose Loc-FloDiff 180180M
Loc-FloDiff (F3) F3Loc Loc-FloDiff 180180M
@@
TABLE III: Error Comparison of Localization Methods
@@
IV-A Main Experiments
Setup
We conduct our experiments in the Gibson [33] simulator. The testing scenes consist of 50 Gibson scenes and 50 HM3D [22] scenes, each containing 10 episodes. The episodes for Gibson scenes are sourced from FloNa [14], while those for HM3D scenes are newly collected under the constraint that the Euclidean distance between the start and target points exceeds 10mm to ensure sufficient task difficulty. The navigation is considered successful only if the agent arrives within a distance of $\tau_{d}$ from the target, within the maximum step limit, and with the total number of collisions below the threshold $\tau_{c}$. The agent responds to collisions by rotating 45 degrees clockwise and then re-predicting its future action sequence.
Baselines
We set eight baselines, with their specific components outlined in Tab. II. Loc-A* (GT): The agent follows a path planned by the A* algorithm, with GT pose used for localization. DD-PPO [32] (GT): This baseline is a reinforcement learning–based agent that navigates using only visual observations, with GT pose for localization. Map-A* (GT): This baseline first uses the same depth estimation model [13] as GlocDiff to convert RGB observations into depth. The depth is then transformed into a point cloud, which is projected to form a bird’s-eye-view (BEV) occupancy grid map. Navigation is carried out by planning a path with the A* algorithm over this occupancy map, using ground truth (GT) pose for localization. Both Loc-FloDiff (F3) and GlocDiff (F3) employ F3Loc [3], an end-to-end floor plan localization model, as their localization module. Loc-FloDiff (GT) and GlocDiff (GT), in contrast, rely on GT pose for localization. In GlocDiff (VO), a VO algorithm [16] is used to track the agent’s pose, serving as the localization module. For evaluation, we adopt Success Rate (SR) and Success Weighted by Path Length (SPL) as our metrics. We evaluate all baselines under the distance threshold $\tau_{d}$ values of $0.25$, as well as the collision threshold $\tau_{c}$ values of $10$, $30$, $50$, and $\infty$.
Figure 4: Trajectory visualization. The visualization of trajectories for different models across six episodes.
Figure 5: Trajectory visualization of GlocDiff (GT) under three distinct global guidance conditions for the same observation.
Quantitative Results
Tab. I presents the qualitative results. Below, we discuss several key observations:
  • DD-PPO (GT) and Map-A* (GT) consistently underperform GlocDiff (GT) across all settings. While Map-A* (GT) achieves slightly higher SR than Loc-FloDiff (GT) at $\tau_{c}=10$ and $30$, its SPL is lower in all cases. This suggests that incorporating a floor plan significantly improves navigation efficiency. Notably, the poor performance of DD-PPO (GT) stems from its lack of explicit obstacle-avoidance capability, combined with our collision-sensitive evaluation—unlike the setup in [32], which allows the agent to slide along obstacles.
  • Despite being trained on a significantly smaller dataset, GlocDiff consistently outperforms Loc-FloDiff by a notable margin in both success rate and SPL, regardless of whether localization is based on GT positions or F3 Loc [3]. Moreover, when $\tau_{c}\geq 30$, GlocDiff (GT) attains a success rate above $90\%$ and an SPL exceeding about $70\%$, strongly indicating that our model effectively utilizes floor plan information and exhibits robust planning capability. These outcomes validate the effectiveness and high data efficiency of our approach. Additionally, when evaluated on HM3D scenes — which differ from the Gibson scenes used in training — GlocDiff exhibits a smaller performance decline than Loc-FloDiff. This indicates that extracting both shortest-path guidance and depth cues from floor plans and RGB observations generalizes more robustly across varied environments compared to methods that rely primarily on learned image features.
  • When the localization results shift from GT poses to predictions from the F3Loc module, Loc-FloDiff experiences a significant decline in performance, with the SR dropping by $32.00\%$ to $41.8\%$ and the SPL decreasing by $29.06\%$ to $38.62\%$. In contrast, the performance of GlocDiff shows a more moderate decline, with the SR falling by $11.20\%$ to $34.60\%$ and the SPL decreasing by $17.59\%$ to $28.77\%$. This highlights that our model exhibits greater robustness to localization noise.
  • The SR of GlocDiff (VO) is higher than that of GlocDiff (F3), with the difference being particularly pronounced under stricter collision constraints (i.e., lower $\tau_{c}$). This suggests that visual odometry (VO) serves as a more suitable localization module for floor plan navigation compared to F3Loc. Since VO can continuously track the agent’s pose, it remains relatively stable for indoor navigation tasks, despite cumulative errors. In contrast, F3Loc maps observations to poses, introducing room-level errors, which is unacceptable for floor plan navigation. We further evaluate the localization accuracy of F3 and VO. Specifically, we compute the mean and median errors between predicted and ground-truth waypoints across fifty sampled trajectories. As shown in Tab. III, VO achieves higher average accuracy than F3Loc. Additionally, F3Loc produces occasional outlier predictions, indicating abrupt jumps to distant locations in the environment.
TABLE IV: Ablation Study on Directional and Depth Cues.
Model w/o depth w/o path GlocDiff (GT)
SR SPL SR SPL SR SPL
$\tau_{c}$ $10$ $64.20$ $47.22$ $44.60$ $38.44$ $79.60$ $62.62$
$30$ $81.00$ $55.98$ $60.60$ $44.09$ $92.80$ $70.38$
$50$ $85.20$ $58.07$ $67.00$ $48.48$ $94.20$ $70.98$
$\infty$ $92.20$ $60.15$ $73.20$ $50.51$ $97.60$ $72.36$
Qualitative Results
Fig. 4 presents qualitative trajectory visualizations of Map-A* (GT), DDPPO (GT), LocFlodiff (GT), and GlocDiff (GT) across six distinct scenarios. We further visualize five trajectories generated by GlocDiff (GT) under three different global guidance inputs for the same observation, as shown in Fig. 5. When the global guidance does not intersect with obstacles (left and right cases), GlocDiff (GT) consistently follows it closely across multiple runs, showing high confidence. In contrast, when the global guidance poses a collision risk with obstacles (middle case), GlocDiff (GT) places greater reliance on local guidance to navigate around them.
IV-B Ablation Study
Ablation Study on Local and Global Guidances
We further evaluate the effects of different guidances on the policy. To assess the impact of the local guidance, we use an EfficientNet to directly encode the RGB observations, obtaining RGB context to replace the original depth context, while keeping other design aspects consistent with those of GlocDiff. The resulting ablation model is denoted as w/o depth. To analyze the influence of the global guidance, we eliminate the global planner and directly utilize the target point to replace the shortest path, while also keeping other aspects consistent with GlocDiff. This modification results in another ablation model, referred to as w/o path. The two models described above are both trained for $20$ epochs and reach convergence.
Figure 6: Performance of different models on the test episodes in Mobridge (left) and Brentsville (right). The left column illustrates the traversed trajectory, while the right column showcases the diverse actions generated by different models based on the same observation.
Figure 7: Three real-world test episodes. The left columns present the top-down view of successful episodes, with $1$ and 55 indicating the start and goal positions, respectively. The right columns show the AGV’s first-person RGB observations at the corresponding positions.
Quantitative results. As outlined in Tab. IV, GlocDiff (GT) achieves the best performance across all cases. Two key conclusions can be drawn. First, the removal of either the local guidance or the global guidance significantly degrades the model’s performance. This indicates that both global and local guidance contribute to enhancing the robot’s ability to navigate safely in complex environments. Second, w/o depth model outperforms w/o path model across SR and SPL metrics. This suggests that explicit global planning based on the floor plan provides more valuable information for efficient navigation than the local depth information derived from observations.
TABLE V: Performance Comparison of w/o Perturbation and GlocDiff
Model w/o perturbation GlocDiff
SR SPL SR SPL
$u_{1}(m)$ $\tau_{c}$ $10$ $61.00$ $42.08$ $73.00$ $53.95$
$30$ $73.00$ $57.15$ $86.00$ $61.86$
$50$ $80.00$ $61.79$ $92.00$ $64.84$
$\infty$ $86.00$ $63.41$ $96.00$ $65.63$
$u_{2}(m)$ $\tau_{c}$ $10$ $42.00$ $29.90$ $65.00$ $44.40$
$30$ $52.00$ $34.41$ $69.00$ $52.07$
$50$ $60.00$ $37.73$ $77.00$ $57.23$
$\infty$ $77.00$ $41.61$ $86.00$ $61.14$
Qualitative results. We select two episodes, each from an unseen scene, and visualize the results in Fig. 6. When the agent navigates to the same goal, different models generate distinct actions due to their varying capabilities. The w/o depth model encounters collisions with walls and furniture in both cases, indicating the role of local guidance in generating collision-free navigation actions. In Mobridge, the w/o path model generates incorrect actions at the intersection due to the lack of guidance from the optimal path.
Ablation Study on GT Perturbation
We retrain our model removing the GT perturbation trick mentioned in Sec. III-E and refer to this revised version as w/o perturbation. To evaluate the robustness of GlocDiff and w/o perturbation to localization noise, we conduct a set of experiments across ten unseen environments. We introduce two levels of noise, $u_{1}(m)\sim\mathcal{U}(0.0,0.5)$ and $u_{2}(m)\sim\mathcal{U}(0.0,1.0)$, to the GT localization results and evaluate the performance of GlocDiff and w/o perturbation under both conditions. As shown in Tab. V, under noisy ground-truth localization, GlocDiff consistently outperforms w/o perturbation in most scenarios. This indicates that incorporating noise during training significantly enhances the model’s robustness to localization errors.
Ablation Study on Prediction Horizon